The following data represent the beginning inventory and, in order of occurrence, the purchases and sales of Blumen, Inc., for an operating period.

Units

Unit Cost

Total Cost

Units Sold

Beginning Inventory

30

$14

$ 420

Sale No. 1

20

Purchase No. 1

50

20

1,000

Sale No. 2

40

Purchase No. 2

20

22

440

Totals

100

$1,860

72


Assuming Blumen, Inc. uses LIFO periodic inventory procedures, the ending inventory cost is:

A.

$620

B.

$560

C.

$880

D.

$840
Answer #1

1) Option A is correct.

Working:

Total units (Beginning + purchased) = 100

Total units sold = 60

Ending inventory units = 100 - 60 = 40

Ending inventory cost LIFO (periodic) = 30 units * $14 + 10 units * 20 = $620

Therefore, option A is correct.

Note:

LIFO means last in first out so sales made out of recent purchases and Ending inventory is from Beginning inventory 30 units and purchases NO.1 10 units